# Learning Quadruped Locomotion Using Differentiable Simulation

### v2

@@ V-C On the Importance of Non-differentiable Terminal Penalty
This subsection highlights one important benefit of Reinforcement Learning (RL) compared to differentiable simulation: RL can significantly enhance its robustness by directly optimizing through non-differentiable rewards or penalties. Specifically, we use an non-differentiable value $p=200$ to penalize the robot (Eq. 7) when the robot experiences termination during training, e.g., falling on the ground.
$r(\mathbf{x}_{t},\mathbf{u}_{t})=\begin{cases}-l(\mathbf{x}_{t},\mathbf{u}_{t})-p&\text{if termination}\\ -l(\mathbf{x}_{t},\mathbf{u}_{t})&\text{otherwise}\end{cases}$ (8)
@@
Fig. 6 shows a study of using non-differentiable terminal penalty for both RL and differentiable simulation. The results show that adding a final penalty can greatly affect how well reinforcement learning (RL) works. Without a penalty, RL might get trapped in a local minimum. However, with a large penalty at the end, RL can achieve better task reward as well as more robust control performance. This is because RL optimizes a discounted return, which estimates “how good” it is to be in a given state. RL uses a state-value function to encode this information
$V_{\pi}(s)=\mathbb{E}[G|S_{0}=s]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|S_{0}=s\right],$ (9)
@@
On the other hand, a terminal penalty has no impact on differentiable simulation since the gradient of a constant value is equal to zero. As a result, differentiable simulation requires well-defined continuous functions, e.g., a potential function or control barrier functions for robust control.
@@
V-D Real World Experiment
Figure 7: Real-world deployment using a Mini Cheetah. All the experiments are conducted using one single blind policy.
We demonstrate the performance of our policy in the real world using Mini Cheetah. Fig. 7 shows several snapshots of the robot’s behavior using different gait patterns. We trained a blind policy in simulation using 64 robots and then transfer the policy directly to the real world without fine-tuning. The robot can walk forward and backward with different gait patterns and frequencies. Moreover, the policy proved robust, enabling the robot to manage certain disruptions, such as unexpected forces applied to its body and locomotion on deformable terrain.
VI Limitations
A fundamental advantage of reinforcement learning (zeroth-order gradient optimization) lies in its optimization objective: RL can directly optimize non-differentiable rewards, including task-level rewards [32]. Conversely, differential simulation relies on information to be backpropagated from its differentiable loss function, making it challenging to explore novel solutions guided by task-level objectives, e.g., how to make survive differentiable?
Our system design requires the specification of foot position using heuristic methods, which is a limitation compared to RL. RL’s flexibility allows for more robust performance enhancements, such as implementing a simple termination penalty to discourage falling behaviors or employing constant survival rewards to promote longevity.
VII Conclusion
In conclusion, we propose a novel framework leveraging the advantages of differentiable simulators for the complex task of quadruped locomotion in the real world. Our framework overcomes the challenges of contact-rich environments for differentiable simulators by integrating a simplified rigid-body model with a non-differentiable model for whole-body dynamics. In this setting, the learned locomotion policy is updated by directly backpropagating the differentiable loss function through the simplified model while the state of the robot is aligned using the more precise whole-body dynamics model. This framework not only introduces the advantages of faster and more stable training convergence of differential simulators to legged robots but also ensures the accurate simulation of complex contact forces. Notably, our experiments demonstrate that the proposed framework can train a policy to walk in minutes using only one simulation environment, a setting in which model-free RL falls short. Furthermore, the policy, aligned with an accurate simulator, seamlessly transfers to the real world in a zero-shot fashion.
An interesting future application of our simulator alignment is the task of learning directly on the hardware, where the locomotion policy is continually updated using the differentiable model in an online fashion while aligning robot states directly with real-world states. Beyond this, the proven training stability and fast training convergence of our differential simulator framework holds the promise for addressing the high-sample inefficiency in tasks like vision-based policy for navigation. In summary, our work bridges the gap between differentiable simulators and real-world legged locomotion, opening avenues for numerous future applications.
VIII Acknowledgments
This work was supported by the European Union’s Horizon Europe Research and Innovation Programme under grant agreement No. 101120732 (AUTOASSESS) and the European Research Council (ERC) under grant agreement No. 864042 (AGILEFLIGHT). The authors especially thank Yuang Zhang for his help with the software development, Seungwoo Hong for the useful discussion about rigid-body dynamics and nonlinear model predictive control, and Nico Messikommer for the valuable feedback. Also, the authors thank Se Hwan Jeon, Ho Jae Lee, Matthew Chignoli, Steve Heim, Elijah Stanger-Jones, and Charles Khazoom for their helpful support.

### v3

@@ Appendix 4 On the Importance of Non-differentiable Terminal Penalty
We highlight one important benefit of RL compared to differentiable simulation: RL can significantly enhance its robustness by directly optimizing through non-differentiable rewards or penalties. Specifically, we use a non-differentiable value $p=200$ to penalize the robot when the robot experiences termination during training, e.g., falling on the ground or lifting its legs above its body.
$r(\mathbf{x}_{t},\mathbf{u}_{t})=\begin{cases}-l(\mathbf{x}_{t},\mathbf{u}_{t})-p&\text{if termination}\\ -l(\mathbf{x}_{t},\mathbf{u}_{t})&\text{otherwise}.\end{cases}$
@@
Fig. 6 shows a study of using non-differentiable terminal penalty for both RL and differentiable simulation. The results show that adding a final penalty can greatly affect how well RL works. Without a penalty, RL might get trapped in a local minimum. However, with a large penalty at the end, RL can achieve better task rewards as well as more robust control performance. This is because RL optimizes a discounted return, which estimates “how good” it is to be in a given state. RL uses a state-value function to encode this information
$V_{\pi}(s)=\mathbb{E}[G|S_{0}=s]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|S_{0}=s\right].$
@@
On the other hand, a terminal penalty has no impact on differentiable simulation since the gradient of a constant value is equal to zero, and we do not leverage a state value function. As a result, differentiable simulation requires well-defined continuous functions, e.g., a potential function or control barrier functions for robust control.
@@
Appendix 5 Ablation Study
We conducted an ablation study to investigate the significance of the proposed state alignment mechanism. In our implementation, we use the state from IsaacGym to align the robot state within our simplified rigid-body differentiable simulation. To assess the impact of state alignment, we performed experiments where we removed the state alignment and compared the resulting learning curves to those obtained with state alignment. The results are presented in Figure 7. The results indicate that without state alignment, the robot fails to learn any useful walking skills.
Figure 7: Ablation study for state alignment using IsaacGym.
Appendix 6 Hyperparameters
Parameter Value
learning rate 0.001
discount factor $\gamma$ 0.95
GAE-$\lambda$ 0.95
learning epoch 10
policy network MLP [256, 256]
value network MLP [256, 256]
clip range 0.2
entropy coefficient 0.002
number of epoch 10
Table 3: PPO hyperparameters.
Parameter Value
learning rate 0.001
policy network MLP [256, 256]
gradient decay factor $\alpha$ 0.9
Table 4: Hyperparameters for policy training using differentiable simulation.